A new virus is released on the internet; the administrator of a department's Local Area Network ( LAN) is given five minutes by a manager to estimate the impact. The administrator samples 15 of the PCs connected to the LAN, and finds that 9 are infected; use proportion to estimate the number of infected PCs if there are a total of 202 PCs connected to the LAN. : A new virus is released on the internet; the administrator of a department's Local Area Network ( LAN) is given five minutes by a manager to estimate the impact. The administrator samples 15 of the PCs connected to the LAN, and finds that 9 are infected; use proportion to estimate the number of infected PCs if there are a total of 202 PCs connected to the LAN.
Answer by jojo14344(879)

If to be based on 15 sample PC which 9 are infected, it follows:
9/15 ---------> cross(9)3/cross(15)5=3/5
For every 5 PC, 3 are infected by proportion.
(3/5)100=60% of the PC are with virus.
So if there are 202 in total:
(202)(60/100)=202*0.60=121.2
Conclusion: Estimated infected about 121 PCs
Thank you,
Jojo